Currently, Nora is 3 times as old as Damon. In 8 years, she will only be 2 times as old as Damon. How old will Damon be in 8 years?



1
Expert's answer
2021-06-28T04:04:21-0400

Let Nora be "x" years old and let Damon be "y" years old.

Nora is 3 times as old as Damon: "x=3y" .

In 8 years, Nora will only be 2 times as old as Damon: "x+8=2(y+8)"


"\begin{cases}
x=3y
\\
x+8=2y+16
\end{cases}\qquad \begin{cases}
x=3y
\\
3y+8=2y+16
\end{cases}\qquad \begin{cases}
x=24
\\
y=8
\end{cases}"


Now Damon is "8" years old. In "8" years Damon’s age will be "8+8=16" .


Answer: "16" years.